OpenStudy (anonymous):

Find an equation of the line tangent to the graph of y=-4/(x^4) at the point (2,-4/16).

OpenStudy (anonymous):

first find derivative of -4/(x^4)

OpenStudy (anonymous):

16x^-5

OpenStudy (anonymous):

same as -4^(-4) use this rule a^n -> n*a^(n-1)

OpenStudy (anonymous):

Then would that be -1/256?

OpenStudy (anonymous):

16x^(-5) plug in 2 16*(2)^(-5)

OpenStudy (anonymous):

16*(2)^(-5)=.5 that is your m

OpenStudy (anonymous):

then do I plug in the y and x with the slope value to find b?

OpenStudy (anonymous):

you are given (2,-4/16). using y=mx+b -4/16=.5*2+b and sfb -1.25 y=.5*x-1.25

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

So the final answer is y=.5x-1.25, right?!

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (anonymous):

Yes